OpenStudy (anonymous):

A statement Sn about the positive integers is given. Write statements S1 S2 and S3 simplifying Sk+1 completely. Sn : 1+ 4 +7 +......+(3n-2)= n(3n-1)/2

OpenStudy (anonymous):

Sn:1+4+7+....(3n-2)=n(3n-1)/2 S1:1=1(3*1-1)/2=1 S2: 1+4=2(3*2-1)/2 or5=5 S3: 1+4+7=3(3*3-1)/2 or 12=3*8/2=12 Assume Sk is true for all positive integers. 1+4+7+.....+(3k-2)=k(3k-1)/2 adding both sides 3(k+1)-2 1+4+7+...+(3k-2)+{3(k+1)-2}=k(3k-1)/2 +3(k+1)-2 =(3k^2-k+6k+6-4)/2 =(3k^2+5k+2)/2 =(3k^2+3k+2k+2)/2 ={3k(k+1)+2(k+1)}/2 =(k+1)(3k+2)/2 =(k+1){3(k+1)-1}/2 or Sk+1=(k+1){3(k+1)-1}/2

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

yw